### v2

# Vision Beyond Boundaries: An Initial Design Space of Domain-specific Large Vision Models in Human-robot Interaction

Abstract.
The emergence of Large Vision Models (LVMs) is following in the footsteps of the recent prosperity of Large Language Models (LLMs) in following years. However, there’s a noticeable gap in structured research applying LVMs to Human-Robot Interaction (HRI), despite extensive evidence supporting the efficacy of vision models in enhancing interactions between humans and robots. Recognizing the vast and anticipated potential, we introduce an initial design space that incorporates domain-specific LVMs, chosen for their superior performance over normal models. We delve into three primary dimensions: HRI contexts, vision-based tasks, and specific domains. The empirical validation was implemented among 15 experts across six evaluated metrics, showcasing the primary efficacy in relevant decision-making scenarios. We explore the process of ideation and potential application scenarios, envisioning this design space as a foundational guideline for future HRI system design, emphasizing accurate domain alignment and model selection.
1. Introduction
The rapid advancement of robotics and human-computer interaction (HCI) has significantly progressed human–robot interaction (HRI), impacting daily life. Key research focuses on understanding interactivity and social behavior between robots and humans (Dautenhahn, 2007), aiming to enable robots to predict human intent and efficiently complete tasks. Literature (Murphy et al., 2010; Kosuge and Hirata, 2004; Sheridan, 2016) highlights HRI’s core aspects: human supervisory control of robots, cooperative task execution, and achieving social goals through interaction. Visual data is crucial in designing intuitive HRI contexts, including gestural interaction (Fan et al., 2022a), object segmentation (Kim et al., 2010), and video tracking (Menezes et al., 2003). Over the past decade, computer vision has evolved remarkably alongside advancements in deep learning. Deep neural networks have enabled significant breakthroughs in tasks like object detection, image segmentation, and scene reconstruction (Wang et al., 2023). More recently, computer vision has seen progress in visual tracking, video captioning, pose estimation, and innovative content generation driven by the emergence of generative artificial intelligence (GenAI), with impressive applications across various domains. Large language models (LLMs) such as ChatGPT, have excelled in text-based tasks, inspiring the development of large vision models (LVMs) which are now pivotal in advancing vision-based analysis and interpretation (Rawte et al., 2023).
The emergence of LVMs has revolutionized HRI by addressing long-standing challenges in visual perception and interpretation (Wang et al., 2023). Originally developed for image-based tasks like object recognition and scene understanding, LVMs are now applied in robotics, enabling the interpretation of complex scenes and informed decision-making with multimodal input. These models enhance robustness and efficiency, similar to how LLMs have transformed text-based applications (Zhang et al., 2023a). However, unlike LLMs, which perform well across varied scenarios, LVMs require domain-specific training due to discrepancies in visual content between formal and informal platforms. Customized LVMs excel in comprehending the unique and nuanced visual content relevant to particular contexts in specific HRI systems. The landscape of AI, particularly in computer vision, has transformed significantly due to the advent of vision transformers and LVMs. Pioneering models like OpenAI’s CLIP (Radford et al., 2021), Landing AI’s LandingLens 111https://landing.ai/platform/, Google’s ViT (Dosovitskiy et al., 2020), and the SWIN Transformer (Liu et al., 2021) have revolutionized visual data processing. Domain-specific LVMs, such as OpenAI’s GPT-4V 222https://openai.com/research/gpt-4v-system-card and Meta’s SAM (Kirillov et al., 2023), focus on specialized enterprise applications, enabling businesses to customize models for their specific needs. These domain-specific LVMs offer advantages over general models, including reduced training costs, enhanced accuracy, and unprecedented scalability through extensive parameter fine-tuning 333https://www.solulab.com/large-vision-models/#What_are_Large_Vision_Models.
According to Landing AI444https://landing.ai/, the LVM revolution lags behind LLMs by two to three years, but domain-specific LVMs outperform general ones. We anticipate regular advancements from these models, leading to more intelligent HRI systems and seamless robot integration into human society. To our knowledge, this work is the first to explore LVMs’ potential in human–robot interaction, contributing an initial design space and empirical evaluation to guide the use of domain-specific LVMs in future HRI systems.
2. Related Work
2.1. Design Space with HRI
In HCI, a design space is a tool that highlights diverse possibilities for crafting specific artifacts, enriching and facilitating the design process. It evolves iteratively, incorporating new potentials based on insights from each design iteration (MacLean et al., 1991). Numerous studies have explored design spaces in HRI. In 2002, Dautenhahn investigated the design space of social robots, identifying niche spaces for human trust (Dautenhahn, 2002). Woods et al. developed a design space to identify children’s perspectives on robots, highlighting ’Behavioral Intention’ and ’Emotional Expression’ as key factors (Woods, 2006). Walters proposed an empirical framework examining social robots’ appearance and behavior, finding that consistent behavior and appearance are preferred by humans (Walters, 2008). Kalegina et al. conducted a survey on human perception of robot faces, noting that certain facial features significantly impact perceptions of a robot’s capabilities and trustworthiness, influencing its appropriateness for specific tasks or environments (Kalegina et al., 2018).
2.2. LVMs in HRI: status
A considerable amount of research has integrated advanced vision models into HRI systems for specialized tasks (Kang and Han, 2023; Ong et al., 2012; Sheron et al., 2021; Dehghan et al., 2019; Sakurai et al., 2007; Martínez and Vitria, 2001; Xu et al., 2020; Lee et al., 2016; Zhao et al., 2018). Object detection is a prevalent task, with vision models enhancing performance and enabling better robot-human interaction (Hsu et al., 2018; Shieh et al., 2013; Latif et al., 2023; Maiettini et al., 2021; Saputra et al., 2019). Azagra et al. proposed a pipeline for interaction type recognition and target object detection (Azagra et al., 2017). Fang et al. developed a hand-held object detection framework using YOLO (Redmon et al., 2016) and KCF (Henriques et al., 2014) on RGB-D video data (Fang et al., 2018). Yu et al. used the LSTM model (Hochreiter and Schmidhuber, 1997) for video object detection in a space robot-human scenario with gestural interactions (Yu et al., 2022). Visual segmentation, which partitions images and videos, is another common task in HRI research (Chavarria et al., 2013; Ghidary et al., 2001; Fei et al., 2023). Ückermann et al. presented a real-time 3D segmentation algorithm for robot grasping with a Shadow Robot Hand (Ückermann et al., 2013). Fan et al. proposed a gesture segmentation and recognition architecture using deep convolutional neural networks (Fan et al., 2022a). Kim et al. designed a multimodal HRI framework for online object segmentation and gesture recognition (Kim et al., 2010). Besides, Visual tracking is also frequently used in HRI research. My et al. (My and Zell, 2013) and Putro et al. (Putro and Jo, 2018) developed real-time face tracking systems for seamless interaction with moving robots, while Song et al. created a face tracking method using image data in a human-mobile robot environment (Song and Chen, 2004). Other tasks like pose estimation, content generation, and visual recognition are also common in HRI studies. However, there is a notable lack of research integrating well-defined LVMs into human-robot interaction, especially domain-specific LVMs. Given the effectiveness of LLMs in text-based information processing, we posit that domain-specific LVMs could transform vision-based contextual analysis across various industries by leveraging their specialized expertise.
3. The Design Space
In this section, we present the formulation of the proposed design space encapsulating three main dimensions.
3.1. HRI Contexts
HRI has become a crucial field at the intersection of robotics and HCI over the past decades. As robots increasingly coexist with humans in various work environments, creating safe and user-friendly systems is essential. The challenge lies in integrating human behaviors into robots’ decision-making frameworks for close physical interactions. With AI technology advancing, HRI research focuses on ensuring safe physical interactions and fostering socially appropriate interactions sensitive to cultural nuances. The goal is to develop robots capable of intuitive communication with humans using speech, gestures, and facial expressions. HRI extends to multiple domains, including industrial automation, medical assistance, and personal companionship. Various studies have categorized different contexts of interactions between humans and robots, yielding similar outcomes based on distinct underlying principles (Schulz et al., 2018; Bütepage and Kragic, 2017; Sheridan, 2016). In our paper, we propose a new and concise categorization of HRI contexts tailored for future LVM deployment (see Figure 1).
Human-initiated: In this context, human supervisory control over robots is initiated for executing routine tasks. It is anticipated that humans will make decisions about the manner and timing of actions, subsequently directing, instructing, or commanding the robots to carry out these actions and provide relevant feedback. Under this framework, robots are considered passive entities, springing into action only upon receiving explicit control signals from human operators.
Robot-proactive: In this context, robots proactively determine the appropriate ways and timing of actions, and also guide humans on their following behaviours. The robots autonomously identify the necessary actions to progress to the subsequent steps required for routine tasks. Concurrently, they independently monitor the situation and observe human actions. Should the predefined goal not be achieved within a set time frame, the robots will then commit an action to ensure the accomplishment of the objective.
Neutral: In this context, both humans and robots proactively initiate actions, collaboratively deciding on the how and when of task execution to achieve mutual goals. Actions required for progressing to the next sub-task are determined simultaneously by both parties. Robots select actions as they would in a robot-proactive context, but they also inform humans about their anticipated contributions, based on the robot’s task plan. A key application of this approach is in social interactions between humans and robots, where robotic devices are designed to entertain, educate, comfort, and assist specific human users.
3.2. Vision-based tasks
The second dimension identified pertains to vision-based tasks, developed by synthesizing insights from published works on visual engagement in HRI and current categorizations of computer vision tasks (Voulodimos et al., 2018) (Figure 1).
Visual Detection: At present, this is the most prevalent task designed in contemporary HRI research employing vision models, such as object detection (Hsu et al., 2018; Shieh et al., 2013) and face detection (Lee et al., 2016; Putro and Jo, 2018). In the majority of cases, the visual detection tasks have significantly enhanced the robustness of interaction with robots.
Visual Classification: As a fundamental task that has evolved alongside deep learning, this area is anticipated to bring significant advantages to HRI systems, particularly in instances like human/robot intention classification (Ong et al., 2012) and object classification in robot teaching (Dehghan et al., 2019).
Pose Estimation: This task is often integrated with visual tracking, enabling robots to proficiently perform 3D estimation of either humans or objects (Xu et al., 2020).
Visual Captioning: As an emerging vision task, it is expected to bring more convenience where robots provide education, or social companionship (Kang and Han, 2023).
Visual Info Generation: With the swift advancements in GenAI, this task has demonstrated its high capability in the field of robotics in recent years. In HRI, we foresee its more utilization, building on the current successes in areas like motion generation (Zhao et al., 2018) and location generation (Saputra et al., 2019).
3.3. Specific domains
Compared to generic LVMs, domain-specific models require only 10% to 30% of the labeled data and produce significantly fewer errors. This advancement reduces computational demands and improves performance. We identified eight domains (Figure 1) expected to encompass most current and prospective HRI application areas, based on insights from published works and research trends.
Healthcare: HRI is revolutionizing health such as robotic-assisted surgery, where precision and reduced invasiveness are critical. Robots like exoskeletons are also being deployed for rehabilitation, providing physical therapy with adaptive routines. The superiority of LVMs in detecting anomalies in medication-related imaging and forecasting disease progression will enhancing the decision-making.
Automotive: LVMs are central in advancing the autonomous vehicles. The real-time object and pedestrian detection, will enable safer navigation and interaction with the environment. Furthermore, LVMs enhance the design of ergonomic vehicle interiors, thereby elevating user experience and safety in driver-assist systems.
Manufacturing: LVMs enable robots to execute precise quality inspections, detect defects, and ensure accurate assembly, collaborating with human operators. In addition, the process monitoring and maintenance significantly improve the efficiency and workplace safety.
Entertainment: LVMs in HRI are anticipated to vastly improve entertainment experiences through their prowess in generating visual content, including enhanced video and gaming experiences. Moreover, the capacity for analyzing visual data, like predicting human intentions, is set to further enrich the quality of interactive experiences between humans and robots.
Security: In critical situations, robots are enabled to perform continuous surveillance, recognize suspicious activities, and identify potential threats with desirable accuracy. Also, utilizing advanced facial recognition and behavior analysis from LVMs to ensure human safety and security publicly.
Agriculture: LVMs enable agricultural robots (such as drones) to accurately identify and classify crops, assess plant health, predict incoming yields, and detect pests or diseases, optimizing the cultivating and harvesting.
Education: This domain has been ubiquitously probed in HRI. With the aid of LVMs, educational robots are to precisely recognize and respond to manifold gestures and expressions from humans, facilitating more engaging and personalized teaching performance. Some vision-related techniques such as virtual reality/augmented reality are able to make learning more immersive and accessible.
Social Interaction: Social robotics is a crucial facet of HRI. In social settings, LVMs enable robots to interpret and respond to human emotions and social cues, fostering more natural and empathetic communication and engagement. A wide-used case is eldercare, where the social robots can provide companionship and emotional support, tailored to individual needs.
4. Evaluation
4.1. Study Design
Exploring the design space’s dimensions provides valuable insights into LVM diversity for HRI researchers and practitioners. However, it doesn’t reveal how these variations affect individual responses. To address this, we conducted an expert evaluation to gather empirical evidence for developing HRI systems that evoke specific reactions to LVMs. We engaged 15 participants aged 24 to 42 (7 male, 6 female, 2 prefer not to say, M = 30.33, SD = 4.66) through email invitations and personal outreach. Participants had several years of experience in relevant areas (see Figure 2.A), enhancing the study’s reliability. After a brief introduction to the design space, we used a questionnaire to explore experts’ perspectives on its three dimensions and the design space itself. This study employed a questionnaire-based evaluation and quantitative analysis to measure five perceived metrics from existing literature (Gries, 2004; Kalegina et al., 2018; Baartmans et al., 2020; Haggerty et al., 2011). The questionnaire included demographic information, an overview of the design space (see Figure 1), and targeted questions to gauge perceptions of likeability, trustworthiness, usefulness, intent to use, and comprehensiveness, using a 7-point Likert scale. The questions are detailed in the appendix.
4.2. Data Analysis
On average, participants completed each trial in about 15 minutes, including a brief interview session to share their opinions. All participants successfully finished the study and received a small gift as appreciation. Before conducting statistical analysis, normality tests confirmed that all metrics adhered to a normal distribution (). We used one-way ANOVA with repeated measures () for the five metrics, wherein, each dimension and the design space received satisfactory ratings across all metrics (Figure 2.B). Notably, the HRI context dimension achieved the highest ratings in every metric, while the vision-based tasks dimension received the lowest.
Likeability. Participants showed an overall preference favoring the HRI contexts dimension as the most likable, closely followed by the design space itself and then the specific domains dimension. However, the differences in likeability ratings did not reach statistical significance.
Trustworthiness. In terms of trustworthiness, perceptions varied slightly from likeability. Here, the specific domains dimension was considered second only to the HRI contexts dimension, though the gap was marginal. Again, these differences were not statistically significant.
Usefulness. The vision-based tasks dimension was seen as the least motivational for future research, distinctly lagging behind the other dimensions. This assessment was supported by a significant statistical difference, as indicated by a repeated measures ANOVA (), highlighting the variance in perceived usefulness across dimensions.
Intent to Use. The variance in the intention for potential adoption was the smallest across all dimensions with the design space itself when compared to other five evaluated metrics. The vision-based tasks dimension was rated only slightly falling behind other dimensions in this metric, although no significance was detected.
Comprehensiveness. Within the whole study, the HRI contexts dimension received the highest rating for comprehensiveness, nonetheless, the vision-based tasks dimension obtained the lowest rating. Statistical significance was affirmed by the repeated measures ANOVA on the comprehensiveness scales with .
4.3. Summary of Qualitative Feedback
The majority of participants expressed positivity of the design space and its three dimensions. One participant, whose research interests align closely with the proposed topic, mentioned, ”The design space you’ve introduced seems like it will be a useful tool and guide for my future research endeavors to a significant extent, as it clearly lists the essential components.” In terms of the three dimensions, another participant appreciated the strategy of outlining specific domains of focus before diving into practical research challenges brought by LVMs: ”This approach of summarizing key domains before tackling practical research problems can streamline the process of identifying the precise area of interest for applying the vision model. It not only saves time but also leads to enhanced outcomes.” Moreover, a participant with years of experience in computer vision and robotics commented on the utility of the HRI contextual categorization, stating, ”This categorization of contexts will aid in laying down the essential structure for the interactive system I plan to develop in the future, especially since my prior experience in this area is limited.”
5. Discussion
In this section, we delve into various facets of the proposed design space.
It serves as a preliminary tool for designers and researchers. In this study, we formulated an initial design space, featuring three dimensions tailored for users to develop future HRI systems using domain-specific LVMs. Our validation process included an empirical testing by experts and the demonstration of its foundational utility across six distinct metrics. Notably, the dimension focusing on HRI contexts received the highest appreciation, whereas the vision-based tasks dimension was evaluated as the least effective according to our metrics. This discrepancy might stem from experts’ perception that, despite our efforts to be exhaustive, the list of vision tasks might not fully encapsulate the breadth of existing research in this area. Given the fast-paced advancements in computer vision, it’s plausible that more innovative and effective vision tasks will emerge. Nevertheless, we believe that the interaction contexts between humans and robots can be effectively categorized within the three classes we proposed, based on the current state of knowledge. The overall design space received satisfactory appraisal in all metrics, implying that its fundamental utility was primarily endorsed by experts.
The ideation is based on the existing work and emerging trends. While numerous studies have focused on integrating various vision models in HRI, primarily for enhanced visual detection and robot control, there is a lack of structured approaches in utilizing LVMs for efficient interaction between humans and robots. Given the current surge in LLMs, we anticipate a similar revolutionary trajectory for LVMs in following years. The concept of domain-specific LVMs, first promoted and advocated by Landing AI, has shown superior performance over traditional models. This underscores the need for a foundational guideline that assists in designing future HRI systems, delineating specific contexts, vision tasks, and domains.
The proposed approach offers several advantages over current models. Utilizing domain-specific unlabeled data substantially reduces the reliance on expensive labeled data, making the development of models more economical. This cost reduction could lead to enhanced interactions with robots. Additionally, the need for less data is capable to yield faster training periods, enabling the rapid execution and implementation of LVMs in real-world HRI applications especially in wearable and social robotics. Furthermore, by selecting datasets that are specialized for specific domains, the models are expected to achieve significantly higher accuracy due to a deeper understanding of the intricacies unique to different areas.
Using the design space. Our proposed design space serves as an effective tool for preliminarily identifying key elements in human-robot collaborative environments. For instance, the development of a robot arm controlled by human gestures to grasp objects includes: a human-initiated context is established, focusing on tasks such as gesture recognition (visual recognition) and object detection (visual detection). The selection and application of LVMs then follow, tailored to the domains (such as education or entertainment), enhancing visual task performance and facilitating smoother control and interaction. Similarly, in a social companion robot scenario, interaction occurs within a neutral context, employing LVMs specialized for social interaction domain. These models can be utilized for recognizing human emotions and behaviors (visual recognition) and providing empathetic responses (visual captioning). Through our expert evaluation, we are expecting that our design space will aid in informed decision-making in relevant cases.
General challenges. Despite their considerable promise, several challenges need addressing for the broader adoption and practical application. A primary issue is accurately determining the appropriate domain before utilizing specific LVMs to guarantee the precision of visual data, as many HRI studies often overlook domain specification. Data availability may be constrained in some fields due to the difficulty of accessing substantial amounts of domain-specific data, particularly in certain industries. Furthermore, establishing effective interaction methods with robots remains an ongoing challenge in the broader field of HRI research.
Ethical considerations. Another significant factor impacting the successful deployment of the design space is ethic concerns. LVMs require a vast number of database, which can raise a significant factor – data bias and fairness, due to the bias inheriting in the training data. This can lead to inequitable or unethical outcomes, especially in sensitive areas like facial recognition in security and healthcare, where privacy concerns are paramount. Additionally, the interpretability and explainability of these large models are crucial; a lack of understanding of how these models function could compromise transparency. The substantial computational resources required for LVMs pose another significant gap, potentially impeding the successful deployment of an effective HRI system. Finally, ethical issues from HRI’s perspective, such as robots being safe and trustworthy rather than threatening data privacy and human well-being, are expected to be addressed (Etemad-Sajadi et al., 2022).
Limitations of the initial design space. While our design space is formulated on existing HRI research and the evolving trends of large foundational models, certain limitations are inherent. We have identified nine vision-based tasks and eight domains aimed at encompassing the most relevant scenarios. However, the continuous advancement of vision models may give rise to new visual tasks in the future, and HRI could potentially expand into currently not-well-explored domains such as pharmacy and electronics. Moreover, the absence of user studies or expert evaluations in our approach means a lack of practical feedback and insights from professionals in the field, which is crucial for comprehensive validation and enhancement of the design space.
6. Conclusion
In this paper, we present an initial design space focused on the future development of HRI systems, leveraging domain-specific LVMs. Mirroring the recent success of LLMs, we foresee LVMs as transformative agents in vision-based information processing within HRI endeavors. We carried out an empirical validation with 15 expert participants, and all perceived metrics and SUS evaluation pointed towards the promising employment of our proposed methodology. Our advocacy for domain-specific models instead of the normal ones stems from their potential to drive significant enhancements and progress in targeted scenarios, enabling more fluid interactions with robots. We believe that this design space will serve as an accelerator for innovative HRI designs across a variety of domains.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
- (1)
- Azagra et al. (2017) Pablo Azagra, Florian Golemo, Yoan Mollard, Manuel Lopes, Javier Civera, and Ana C Murillo. 2017. A multimodal dataset for object model learning from natural human-robot interaction. In 2017 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, 6134–6141.
- Baartmans et al. (2020) Jeanine MD Baartmans, Francisca JA van Steensel, Lynn Mobach, Tessa AM Lansu, Geraly Bijsterbosch, Iris Verpaalen, Ronald M Rapee, Natasha Magson, Susan M Bögels, Mike Rinck, et al. 2020. Social anxiety and perceptions of likeability by peers in children. British Journal of Developmental Psychology 38, 2 (2020), 319–336.
- Bütepage and Kragic (2017) Judith Bütepage and Danica Kragic. 2017. Human-robot collaboration: from psychology to social robotics. arXiv preprint arXiv:1705.10146 (2017).
- Chavarria et al. (2013) Harold Vasquez Chavarria, Hugo Jair Escalante, and L Enrique Sucar. 2013. Simultaneous segmentation and recognition of hand gestures for human-robot interaction. In 2013 16th International Conference on Advanced Robotics (ICAR). IEEE, 1–6.
- Dautenhahn (2002) Kerstin Dautenhahn. 2002. Design spaces and niche spaces of believable social robots. In Proceedings. 11th ieee international workshop on robot and human interactive communication. IEEE, 192–197.
- Dautenhahn (2007) Kerstin Dautenhahn. 2007. Socially intelligent robots: dimensions of human–robot interaction. Philosophical transactions of the royal society B: Biological sciences 362, 1480 (2007), 679–704.
- Dehghan et al. (2019) Masood Dehghan, Zichen Zhang, Mennatullah Siam, Jun Jin, Laura Petrich, and Martin Jagersand. 2019. Online object and task learning via human robot interaction. In 2019 international conference on robotics and automation (ICRA). IEEE, 2132–2138.
- Dosovitskiy et al. (2020) Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, et al. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv preprint arXiv:2010.11929 (2020).
- Etemad-Sajadi et al. (2022) Reza Etemad-Sajadi, Antonin Soussan, and Théo Schöpfer. 2022. How ethical issues raised by human–robot interaction can impact the intention to use the robot? International journal of social robotics 14, 4 (2022), 1103–1115.
- Fan et al. (2022a) Jinlong Fan, Yang Yue, Yu Wang, Bei Wan, Xudong Li, and Gengpai Hua. 2022a. A continuous gesture segmentation and recognition method for human-robot interaction. In Journal of Physics: Conference Series, Vol. 2213. IOP Publishing, 012039.
- Fan et al. (2022b) Junming Fan, Pai Zheng, and Shufei Li. 2022b. Vision-based holistic scene understanding towards proactive human–robot collaboration. Robotics and Computer-Integrated Manufacturing 75 (2022), 102304.
- Fang et al. (2018) Zhiwen Fang, Junsong Yuan, and Nadia Magnenat-Thalmann. 2018. Understanding human-object interaction in RGB-D videos for human robot interaction. In Proceedings of Computer Graphics International 2018. 163–167.
- Fei et al. (2023) Haolin Fei, Ziwei Wang, Darren Williams, and Andrew Kennedy. 2023. Hybrid Approach for Efficient and Accurate Category-Agnostic Object Detection and Localization with Image Queries in Human-Robot Interaction. In IECON 2023-49th Annual Conference of the IEEE Industrial Electronics Society. IEEE, 1–6.
- Gandarias et al. (2018) Juan M Gandarias, Jesús M Gómez-de Gabriel, and Alfonso J García-Cerezo. 2018. Enhancing perception with tactile object recognition in adaptive grippers for human–robot interaction. Sensors 18, 3 (2018), 692.
- Ghidary et al. (2001) Saeed Shiry Ghidary, Yasushi Nakata, Hiroshi Saito, Motofumi Hattori, and Toshi Takamori. 2001. Multi-modal human robot interaction for map generation. In Proceedings 2001 IEEE/RSJ International Conference on Intelligent Robots and Systems. Expanding the Societal Role of Robotics in the the Next Millennium (Cat. No. 01CH37180), Vol. 4. IEEE, 2246–2251.
- Gries (2004) Matthias Gries. 2004. Methods for evaluating and covering the design space during early design development. Integration 38, 2 (2004), 131–183.
- Haggerty et al. (2011) Jeannie L Haggerty, Marie-Dominique Beaulieu, Raynald Pineault, Frederick Burge, Jean-Frédéric Lévesque, Darcy A Santor, Fatima Bouharaoui, and Christine Beaulieu. 2011. Comprehensiveness of care from the patient perspective: comparison of primary healthcare evaluation instruments. Healthcare policy 7, Spec Issue (2011), 154.
- Henriques et al. (2014) João F Henriques, Rui Caseiro, Pedro Martins, and Jorge Batista. 2014. High-speed tracking with kernelized correlation filters. IEEE transactions on pattern analysis and machine intelligence 37, 3 (2014), 583–596.
- Hochreiter and Schmidhuber (1997) Sepp Hochreiter and Jürgen Schmidhuber. 1997. Long short-term memory. Neural computation 9, 8 (1997), 1735–1780.
- Hsu et al. (2018) Shih-Chung Hsu, Yu-Wen Wang, and Chung-Lin Huang. 2018. Human object identification for human-robot interaction by using fast R-CNN. In 2018 Second IEEE International Conference on Robotic Computing (IRC). IEEE, 201–204.
- Kalegina et al. (2018) Alisa Kalegina, Grace Schroeder, Aidan Allchin, Keara Berlin, and Maya Cakmak. 2018. Characterizing the design space of rendered robot faces. In Proceedings of the 2018 ACM/IEEE International Conference on Human-Robot Interaction. 96–104.
- Kang and Han (2023) Soo-Han Kang and Ji-Hyeong Han. 2023. Video captioning based on both egocentric and exocentric views of robot vision for human-robot interaction. International Journal of Social Robotics 15, 4 (2023), 631–641.
- Kim et al. (2010) Soohwan Kim, Dong Hwan Kim, and Sung-Kee Park. 2010. On-line object segmentation through human-robot interaction. In 2010 IEEE/RSJ International Conference on Intelligent Robots and Systems. IEEE, 1734–1739.
- Kirillov et al. (2023) Alexander Kirillov, Eric Mintun, Nikhila Ravi, Hanzi Mao, Chloe Rolland, Laura Gustafson, Tete Xiao, Spencer Whitehead, Alexander C Berg, Wan-Yen Lo, et al. 2023. Segment anything. arXiv preprint arXiv:2304.02643 (2023).
- Kosuge and Hirata (2004) Kazuhiro Kosuge and Yasuhisa Hirata. 2004. Human-robot interaction. In 2004 IEEE International Conference on Robotics and Biomimetics. IEEE, 8–11.
- Latif et al. (2023) Arsalan Latif, Aimen Mughal, Muhammad Hasan Danish Khan, and Muhammad D Khan. 2023. Human robot Interaction–Object Detection and Distance Measurement Using Kinect V2. In 2023 International Conference on IT and Industrial Technologies (ICIT). IEEE, 1–5.
- Lee et al. (2016) Sang-Seol Lee, Sung-Joon Jang, Jungho Kim, and Byeongho Choi. 2016. A hardware architecture of face detection for human-robot interaction and its implementation. In 2016 IEEE International Conference on Consumer Electronics-Asia (ICCE-Asia). IEEE, 1–2.
- Liu et al. (2021) Ze Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and Baining Guo. 2021. Swin transformer: Hierarchical vision transformer using shifted windows. In Proceedings of the IEEE/CVF international conference on computer vision. 10012–10022.
- MacLean et al. (1991) Allan MacLean, Richard Young, Victoria Bellotti, and Thomas Moran. 1991. Design space analysis: Bridging from theory to practice via design rationale. Proceedings of Esprit 91, 720-730 (1991), 2.
- Maiettini et al. (2021) Elisa Maiettini, Vadim Tikhanoff, and Lorenzo Natale. 2021. Weakly-supervised object detection learning through human-robot interaction. In 2020 IEEE-RAS 20th International Conference on Humanoid Robots (Humanoids). IEEE, 392–399.
- Martínez and Vitria (2001) Aleix M Martínez and Jordi Vitria. 2001. Clustering in image space for place recognition and visual annotations for human-robot interaction. IEEE Transactions on Systems, Man, and Cybernetics, Part B (Cybernetics) 31, 5 (2001), 669–682.
- Menezes et al. (2003) P Menezes, L Brethes, F Lerasle, P Danes, and J Dias. 2003. Visual tracking of silhouettes for human-robot interaction. In Proceedings of The 11th International Conference on Advanced Robotics (ICAR 2003), Coimbra, Portugal. 971–986.
- Molina-Tanco et al. (2005) Luis Molina-Tanco, JP Bandera, Rebeca Marfil, and F Sandoval. 2005. Real-time human motion analysis for human-robot interaction. In 2005 IEEE/RSJ International Conference on Intelligent Robots and Systems. IEEE, 1402–1407.
- Murphy et al. (2010) Robin R Murphy, Tatsuya Nomura, Aude Billard, and Jennifer L Burke. 2010. Human–robot interaction. IEEE robotics & automation magazine 17, 2 (2010), 85–89.
- My and Zell (2013) Vo Duc My and Andreas Zell. 2013. Real time face tracking and pose estimation using an adaptive correlation filter for human-robot interaction. In 2013 European Conference on Mobile Robots. IEEE, 119–124.
- Ong et al. (2012) Kai Siang Ong, Yuan Han Hsu, and Li Chen Fu. 2012. Sensor fusion based human detection and tracking system for human-robot interaction. In 2012 IEEE/RSJ International Conference on Intelligent Robots and Systems. IEEE, 4835–4840.
- Putro and Jo (2018) Muhamad Dwisnanto Putro and Kang-Hyun Jo. 2018. Real-time face tracking for human-robot interaction. In 2018 International Conference on Information and Communication Technology Robotics (ICT-ROBOT). IEEE, 1–4.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
- Rawte et al. (2023) Vipula Rawte, Amit Sheth, and Amitava Das. 2023. A survey of hallucination in large foundation models. arXiv preprint arXiv:2309.05922 (2023).
- Redmon et al. (2016) Joseph Redmon, Santosh Divvala, Ross Girshick, and Ali Farhadi. 2016. You only look once: Unified, real-time object detection. In Proceedings of the IEEE conference on computer vision and pattern recognition. 779–788.
- Sakurai et al. (2007) Shoichiro Sakurai, Eri Sato, and Toru Yamaguchi. 2007. Recognizing pointing behavior using image processing for human-robot interaction. In 2007 IEEE/ASME international conference on advanced intelligent mechatronics. IEEE, 1–6.
- Saputra et al. (2019) Azhar Aulia Saputra, Chin Wei Hong, and Naoyuki Kubota. 2019. Real-time grasp affordance detection of unknown object for robot-human interaction. In 2019 IEEE International Conference on Systems, Man and Cybernetics (SMC). IEEE, 3093–3098.
- Schulz et al. (2018) Ruth Schulz, Philipp Kratzer, and Marc Toussaint. 2018. Preferred interaction styles for human-robot collaboration vary over tasks with different action types. Frontiers in neurorobotics 12 (2018), 36.
- Sheridan (2016) Thomas B Sheridan. 2016. Human–robot interaction: status and challenges. Human factors 58, 4 (2016), 525–532.
- Sheron et al. (2021) PS Febin Sheron, KP Sridhar, S Baskar, and P Mohamed Shakeel. 2021. Projection-dependent input processing for 3D object recognition in human robot interaction systems. Image and Vision Computing 106 (2021), 104089.
- Shieh et al. (2013) Ming-Yuan Shieh, Yen-Hao Chen, Jeng-Han Li, Neng-Sheng Pai, and Juing-Shian Chiou. 2013. Fast object detection for human-robot interaction control. In Proceedings of the 2013 IEEE/SICE International Symposium on System Integration. IEEE, 616–619.
- Song and Chen (2004) Kai-Tai Song and Wen-Jun Chen. 2004. Face recognition and tracking for human-robot interaction. In 2004 IEEE International Conference on Systems, Man and Cybernetics (IEEE Cat. No. 04CH37583), Vol. 3. IEEE, 2877–2882.
- Ückermann et al. (2013) Andre Ückermann, Robert Haschke, and Helge Ritter. 2013. Realtime 3D segmentation for human-robot interaction. In 2013 IEEE/RSJ International Conference on Intelligent Robots and Systems. IEEE, 2136–2143.
- Voulodimos et al. (2018) Athanasios Voulodimos, Nikolaos Doulamis, Anastasios Doulamis, Eftychios Protopapadakis, et al. 2018. Deep learning for computer vision: A brief review. Computational intelligence and neuroscience 2018 (2018).
- Walters (2008) Michael L Walters. 2008. The design space for robot appearance and behaviour for social robot companions. Ph. D. Dissertation.
- Wang et al. (2023) Jiaqi Wang, Zhengliang Liu, Lin Zhao, Zihao Wu, Chong Ma, Sigang Yu, Haixing Dai, Qiushi Yang, Yiheng Liu, Songyao Zhang, et al. 2023. Review of large vision models and visual prompt engineering. Meta-Radiology (2023), 100047.
- Woods (2006) Sarah Woods. 2006. Exploring the design space of robots: Children’s perspectives. Interacting with Computers 18, 6 (2006), 1390–1418.
- Xu et al. (2020) Chengjun Xu, Xinyi Yu, Zhengan Wang, and Linlin Ou. 2020. Multi-view human pose estimation in human-robot interaction. In IECON 2020 The 46th Annual Conference of the IEEE Industrial Electronics Society. IEEE, 4769–4775.
- Yu et al. (2022) Jiahui Yu, Hongwei Gao, Yongquan Chen, Dalin Zhou, Jinguo Liu, and Zhaojie Ju. 2022. Deep object detector with attentional spatiotemporal LSTM for space human–robot interaction. IEEE Transactions on human-machine systems 52, 4 (2022), 784–793.
- Zhang et al. (2023a) Ceng Zhang, Junxin Chen, Jiatong Li, Yanhong Peng, and Zebing Mao. 2023a. Large language models for human-robot interaction: A review. Biomimetic Intelligence and Robotics (2023), 100131.
- Zhang et al. (2023b) Zeyu Zhang, Lexing Zhang, Zaijin Wang, Ziyuan Jiao, Muzhi Han, Yixin Zhu, Song-Chun Zhu, and Hangxin Liu. 2023b. Part-level scene reconstruction affords robot interaction. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, 11178–11185.
- Zhao et al. (2018) Xuan Zhao, Sakmongkon Chumkamon, Shuanda Duan, Juan Rojas, and Jia Pan. 2018. Collaborative human-robot motion generation using LSTM-RNN. In 2018 IEEE-RAS 18th International Conference on Humanoid Robots (Humanoids). IEEE, 1–9.
Appendix A The questions designed for the study
How much do you like this dimension/design space? (Very dislike – Very like)
How trustworthy do you consider this dimension/design space to be? (Very untrustworthy – Very trustworthy)
How useful do you find this dimension/design space in guiding or inspiring future research in related areas? (Very useless – Very useful)
How likely are you to utilize this dimension/design space in designing future HRI systems that incorporate LVMs? (Very unlikely – Very likely)
How comprehensive do you find the dimension/design space in covering the aspects relevant to the future research in related areas? (Very incomprehensive – Very comprehensive)